I need help with all A-D I want to compare and make sure I am doing it right.
Answer #1

A.

Formula for Break Even point (in units)= Fixed cost/Contribution in units

Therefore, BEP(UNITS)= (525000+400000)/31.5*=29319 units

*1577350/50000=$31.55

BEP(Revenue): Product A=29319*140/5= $820932

Product B= 29319*80*4/5= $1876416

B.

Cost Volume Profit Chart

PARTICULARS PRODUCT A PRODUCT B ($)
Sales (Volume) 10000 40000
Sales in $ 1400000 3200000
less: Sales commission @10% (140000) (320000)
NET SALES 1260000 2880000 4140000
less: Variable Cost
Material 375000 600000
Labour 500000 1000000
Manufacturing 1000000 1000000
Overtime 31250 50000
Utilities 1280 5120
TOTAL COST 1907530 2655120 4562650
Contribution (647530) 224880 (422650)
Less: Fixed Cost (925000)
Profit/(Loss) 1347650

C.

If sales Mix changes:

PARTICULARS PRODUCT A PRODUCT B ($)
Sales (Volume) 20000 30000
Sales in $ 2800000 2400000
less: Sales commission @10% (280000) (240000)
NET SALES 2520000 2160000 4680000
less: Variable Cost
Material 750000 450000
Labour 600000 900000
Manufacturing 2000000 750000
Overtime 32500 48750
Utilities 2560 3840
TOTAL COST 3385060 2152590 5537650
Contribution (865050) 7410 (857650)
Less: Fixed Cost (925000)
Profit/(Loss) (1782650)